# Bias and Fairness in Large Language Models: A Survey

### v2

@@ Abstract
Rapid advancements of large language models (LLMs) have enabled the processing, understanding, and generation of human-like text, with increasing integration into systems that touch our social sphere. Despite this success, these models can learn, perpetuate, and amplify harmful social biases. In this paper, we present a comprehensive survey of bias evaluation and mitigation techniques for LLMs. We first consolidate, formalize, and expand notions of social bias and fairness in natural language processing, defining distinct facets of harm and introducing several desiderata to operationalize fairness for LLMs. We then unify the literature by proposing three intuitive taxonomies, two for bias evaluation, namely metrics and datasets, and one for mitigation. Our first taxonomy of metrics for bias evaluation disambiguates the relationship between metrics and evaluation datasets, and organizes metrics by the different levels at which they operate in a model: embeddings, probabilities, and generated text. Our second taxonomy of datasets for bias evaluation categorizes datasets by their structure as counterfactual inputs or prompts, and identifies the targeted harms and social groups; we also release a consolidation of publicly-available datasets for improved access. Our third taxonomy of techniques for bias mitigation classifies methods by their intervention during pre-processing, in-training, intra-processing, and post-processing, with granular subcategories that elucidate research trends. Finally, we identify open problems and challenges for future work. Synthesizing a wide range of recent research, we aim to provide a clear guide of the existing literature that empowers researchers and practitioners to better understand and prevent the propagation of bias in LLMs.
@@ 1 Introduction
Warning: This paper contains explicit statements of offensive or upsetting language.
@@
The rise and rapid advancement of large language models (LLMs) has fundamentally changed language technologies (e.g., Brown et al., 2020; Conneau et al., 2020; Devlin et al., 2019; Lewis et al., 2020; Liu et al., 2019; OpenAI, 2023; Radford et al., 2018; 2019; Raffel et al., 2020). With the ability to generate human-like text, as well as adapt to a wide array of natural language processing (NLP) tasks, the impressive capabilities of these models have initiated a paradigm shift in the development of language models. Instead of training task-specific models on relatively small task-specific datasets, researchers and practitioners can use LLMs as foundation models that can be fine-tuned for particular functions (Bommasani et al., 2021). Even without fine-tuning, foundation models increasingly enable few- or zero-shot capabilities for a wide array of scenarios like classification, question-answering, logical reasoning, fact retrieval, information extraction, and more, with the task described in a natural language prompt to the model and few or no labeled examples (e.g., Brown et al., 2020; Kojima et al., 2022; Liu et al., 2023a; Radford et al., 2019; Wei et al., 2022; Zhao et al., 2021).
@@
Laying behind these successes, however, is the potential to perpetuate harm. Typically trained on an enormous scale of uncurated Internet-based data, LLMs inherit stereotypes, misrepresentations, derogatory and exclusionary language, and other denigrating behaviors that disproportionately affect already-vulnerable and marginalized communities (Bender et al., 2021; Dodge et al., 2021; Sheng et al., 2021b). These harms are forms of "social bias," a subjective and normative term we broadly use to refer to disparate treatment or outcomes between social groups that arise from historical and structural power asymmetries, which we define and discuss in Section 2.111Unless otherwise specified, our use of ”bias” refers to social bias, defined in Definition 7. Though LLMs often reflect existing biases, they can amplify these biases too; in either case, the automated reproduction of injustice can reinforce systems of inequity (Benjamin, 2020). From negative sentiment and toxicity directed towards some social groups, to stereotypical linguistic associations, to lack of recognition of certain language dialects, the presence of biases of LLMs have been well-documented (e.g., Blodgett & O’Connor, 2017; Hutchinson et al., 2020; Mei et al., 2023; Měchura, 2022; Mozafari et al., 2020; Sap et al., 2019; Sheng et al., 2019).
@@
With the growing recognition of the biases embedded in LLMs has emerged an abundance of works proposing techniques to measure or remove social bias, primarily organized by (1) metrics for bias evaluation, (2) datasets for bias evaluation, and (3) techniques for bias mitigation. In this survey, we categorize, summarize, and discuss each of these three areas of research. For each area, we propose an intuitive taxonomy structured around the types of interventions to which a researcher or practitioner has access. Metrics for bias evaluation are organized by the underlying data structure assumed by the metric, which may differ depending on access to the LLM (i.e., can the user access model-assigned token probabilities, or only generated text output?). Datasets are similarly categorized by their structure. Techniques for bias mitigation are organized around the stage of intervention: pre-processing, in-training, intra-processing, and post-processing.
@@ 2.1 Preliminaries
A large language model (LLM) $\mathcal{M}$ parameterized by $\theta$ is a Transformer-based model with an autoregressive, autoencoding, or encoder-decoder architecture that has been trained on a large corpus of hundreds of millions to trillions of tokens. LLMs encompass pre-trained models.
Examples of autoregressive models include GPT (Radford et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and GPT-4 (OpenAI, 2023); autoencoding models include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and XLM-R (Conneau et al., 2020); and encoder-decoder models include BART (Lewis et al., 2020) and T5 (Raffel et al., 2020).
@@
To quantify the performance of an LLM – whether for a downstream task, bias mitigation, or otherwise – an evaluation dataset and metric are typically used. Though benchmark datasets and their associated metrics are often conflated, the evaluation dataset and metric are distinct entities in an evaluation framework, and thus we define a general LLM metric here. In particular, the structure of a dataset may determine which set of metrics are appropriate, but a metric is rarely restricted to a single benchmark dataset. We discuss the relationship between metrics and datasets in more detail in Sections 3 and 4.
@@ 2.2.1 Social Bias and Fairness
Measuring and mitigating social "bias" to ensure "fairness" in NLP systems has featured prominently in recent literature. Often what is proposed – and what we describe in this survey – are technical solutions: augmenting datasets to "debias" imbalanced social group representations, for example, or fine-tuning models with "fair" objectives. Despite the growing emphasis on addressing these issues, bias and fairness research in LLMs often fails to precisely describe the harms of model behaviors: who is harmed, why the behavior is harmful, and how the harm reflects and reinforces social principles or hierarchies (Blodgett et al., 2020). Many approaches, for instance, assume some implicitly desirable criterion (e.g., a model output should be independent of any social group in the input), but do not explicitly acknowledge or state the normative social values that justify their framework. Others lack consistency in their definitions of bias, or do not seriously engage with the relevant power dynamics that perpetuate the underlying harm (Blodgett et al., 2021). Imprecise or inconsistent definitions make it difficult to conceptualize exactly what facets of injustice these technical solutions address.
@@
Here we attempt to disambiguate the types of harms that may emerge from LLMs, building on the definitions in machine learning works by Barocas et al. (2019), Bender et al. (2021), Blodgett et al. (2020), Crawford (2017), Mehrabi et al. (2021), Suresh & Guttag (2021), and Weidinger et al. (2022), and following extensive (socio)linguistic research in this area by Beukeboom & Burgers (2019), Craft et al. (2020), Loudermilk (2015), Maass (1999), and others. Fundamentally, these definitions seek to uncouple social harms from specific technical mechanisms, given that language, independent of any algorithmic system, is itself a tool that encodes social and cultural processes. Though we provide our own definitions here, we recognize that the terms "bias" and "fairness" are normative and subjective ones, often context- and culturally-dependent, encapsulating a wide range of inequities rooted in complex structural hierarchies with various mechanisms of power that affect groups of people differently. Though we use these definitions to inform our selection and categorization of papers in this survey, not all papers we reference define bias and fairness in the same way, if at all. Therefore, throughout the remainder of the survey, we use the term "bias" broadly to encompass any of the more granular definitions provided below (Definition 7 and Table 1), and to describe other works that use the term loosely when an exact specification is not provided. Note that our use of the terms "debiased" or "unbiased" does not mean that bias has been completely removed, but rather refers to the output of a bias mitigation technique, regardless of that technique’s effectiveness, reflecting language commonly used in prior works. Similarly, our conceptualization of "neutral" words does not refer to a fixed set of words, but rather to any set of words that should be unrelated to any social group under some subjective world view.
@@
The primary emphasis of bias evaluation and mitigation efforts for LLMs focus on group notions of fairness, which center on disparities between social groups, following group fairness definitions in the literature (Chouldechova, 2017; Hardt et al., 2016; Kamiran & Calders, 2012). We also discuss individual fairness (Dwork et al., 2012). We provide several definitions that describe our notions of bias and fairness for NLP tasks, which we leverage throughout the remainder of the paper.
@@
We highlight that social groups are often socially constructed, a form of classification with delineations that are not static and may be contested (Hanna et al., 2020). The labeling of groups may grant legitimacy to these boundaries, define relational differences between groups, and reinforce social hierarchies and power imbalances, often with very real and material consequences that can segregate, marginalize, and oppress (Beukeboom & Burgers, 2019; Hanna et al., 2020). The harms experienced by each social group vary greatly, due to distinct historical, structural, and institutional forces of injustice that may operate vastly differently for, say, race and gender, and also apply differently across intersectional identities. However, we also emphasize that evaluating and bringing awareness to disparities requires access to social groups. Thus, under the lens of disparity assessment, and following the direction of recent literature in bias evaluation and mitigation for LLMs, we proceed with this notion of social groups. We now define our notions of fairness and bias, in the context of LLMs.
@@
The taxonomy of bias issues follows those similarly defined by Barocas et al. (2019), Blodgett et al. (2020), and Crawford (2017). Each form of bias described in Table 1 represents a distinct form of mistreatment, but the harms are not necessarily mutually exclusive nor independent; for instance, representational harms can in turn perpetuate allocational harms. We hope that researchers and practitioners can leverage the taxonomy and definitions to describe work in bias mitigation and evaluation with more precise language.
@@
Representational Harms
Perpetuation of denigrating and subordinating attitudes towards a social group
Derogatory language
Pejorative slurs, insults, or other words or phrases that target and denigrate a social group
e.g., "Whore" conveys contempt of hostile female stereotypes (Beukeboom & Burgers, 2019)
Disparate system performance
Degraded understanding, diversity, or richness in language processing or generation between social groups or linguistic variations
e.g., AAE* like "he woke af" is misclassified as not English more often than $\textrm{}^{\dagger}$ equivalents (Blodgett & O’Connor, 2017)
Exclusionary norms
Reinforced normativity of the dominant social group and implicit exclusion or devaluation of other groups
e.g., "Both genders" excludes non-binary gender identities (Bender et al., 2021)
Misrepresentation
An incomplete or non-representative distribution of the sample population generalized to a social group
@@
Stereotyping
Negative, generally immutable abstractions about a labeled social group
e.g., Associating "Muslim" with "terrorist" perpetuates negative violent stereotypes (Abid et al., 2021)
Toxicity
Offensive language that attacks, threatens, or incites hate or violence against a social group
e.g., "I hate Latinos" is disrespectful, hateful, and unreasonable (Dixon et al., 2018)
Allocational Harms
Disparate distribution of resources or opportunities between social groups
Direct discrimination
Disparate treatment due explicitly to membership of a social group
e.g., LLM-aided resume screening may perpetuate inequities in hiring (Ferrara, 2023)
Indirect discrimination
Disparate treatment despite facially neutral consideration towards social groups, due to proxies or other implicit factors
@@ 2.2.2 Bias in NLP Tasks
Language is closely tied to identity, social relations, and power. Language can make concrete the categorization and differentiation of social groups, giving voice to generic or derogatory labels, and linking categories of people to stereotypical, unrepresentative, or overly general characteristics (Beukeboom & Burgers, 2019; Maass, 1999). Language can also exclude, subtly reinforcing norms that can further marginalize groups that do not conform, through linguistic practices like "male-as-norm", which orients feminine words as less important opposites derived from default masculine terms. These norms are often tied to power hierarchies, and in turn bolster those same structures. Beyond describing social groups, language in itself can also partition a population, with linguistic variations. Linguistic profiling, for instance, can discriminate against speakers of a dialect considered non-standard (Baugh, 2000; Loudermilk, 2015). In fact, the determination of which forms of language are considered standard or correct also reinforces social hierarchies that can justify the inferiority of some groups (Blodgett et al., 2020; Craft et al., 2020). Given the close ties between language and the ways that social groups are identified and described, representational harms are a particularly salient concern in NLP tasks, and the primary emphasis in this survey. Of course, representational harms often arise subtly, and thus quantifying them in language, at least for some NLP tasks, differ from standard fairness techniques, which typically apply to classification. We provide a non-exhaustive list of examples of settings where bias may manifest in unique forms, depending on the task:
@@
  • Machine Translation: Machine translators may default to masculine words in the case of ambiguity, a form of an exclusionary norm, such as translating from English "I am happy" to the masculine French form "je suis heureux", as opposed to the feminine form "je suis heureuse(Měchura, 2022).
  • Information Retrieval: Retrieved documents may exhibit similar exclusionary norms as machine translation models, returning more documents with masculine-related concepts than feminine ones, despite using non-gendered queries (Rekabsaz & Schedl, 2020).
@@
  • Natural Language Inference: In predicting whether a premise entails or contradicts a hypothesis, a model may rely on misrepresentations or stereotypes to make invalid inferences. For example, a biased model may infer that "the accountant ate a bagel" entails or contradicts "the man ate a bagel" or "the woman ate a bagel", when the relationship should instead be neutral (Dev et al., 2020).
  • Classification: Toxicity detection models misclassify African-American English tweets as negative more often that those written in Standard American English (Mozafari et al., 2020; Sap et al., 2019).

### v3

@@ Abstract
Rapid advancements of large language models (LLMs) have enabled the processing, understanding, and generation of human-like text, with increasing integration into systems that touch our social sphere. Despite this success, these models can learn, perpetuate, and amplify harmful social biases. In this article, we present a comprehensive survey of bias evaluation and mitigation techniques for LLMs. We first consolidate, formalize, and expand notions of social bias and fairness in natural language processing, defining distinct facets of harm and introducing several desiderata to operationalize fairness for LLMs. We then unify the literature by proposing three intuitive taxonomies, two for bias evaluation, namely metrics and datasets, and one for mitigation. Our first taxonomy of metrics for bias evaluation disambiguates the relationship between metrics and evaluation datasets, and organizes metrics by the different levels at which they operate in a model: embeddings, probabilities, and generated text. Our second taxonomy of datasets for bias evaluation categorizes datasets by their structure as counterfactual inputs or prompts, and identifies the targeted harms and social groups; we also release a consolidation of publicly-available datasets for improved access. Our third taxonomy of techniques for bias mitigation classifies methods by their intervention during pre-processing, in-training, intra-processing, and post-processing, with granular subcategories that elucidate research trends. Finally, we identify open problems and challenges for future work. Synthesizing a wide range of recent research, we aim to provide a clear guide of the existing literature that empowers researchers and practitioners to better understand and prevent the propagation of bias in LLMs.
@@ 1 Introduction
Warning: This article contains explicit statements of offensive or upsetting language.
@@
The rise and rapid advancement of large language models (LLMs) has fundamentally changed language technologies (e.g., Brown et al., 2020; Conneau et al., 2020; Devlin et al., 2019; Lewis et al., 2020; Liu et al., 2019; OpenAI, 2023; Radford et al., 2018, 2019; Raffel et al., 2020). With the ability to generate human-like text, as well as adapt to a wide array of natural language processing (NLP) tasks, the impressive capabilities of these models have initiated a paradigm shift in the development of language models. Instead of training task-specific models on relatively small task-specific datasets, researchers and practitioners can use LLMs as foundation models that can be fine-tuned for particular functions (Bommasani et al., 2021). Even without fine-tuning, foundation models increasingly enable few- or zero-shot capabilities for a wide array of scenarios like classification, question-answering, logical reasoning, fact retrieval, information extraction, and more, with the task described in a natural language prompt to the model and few or no labeled examples (e.g., Brown et al., 2020; Kojima et al., 2022; Liu et al., 2023; Radford et al., 2019; Wei et al., 2022; Zhao et al., 2021).
@@
Laying behind these successes, however, is the potential to perpetuate harm. Typically trained on an enormous scale of uncurated Internet-based data, LLMs inherit stereotypes, misrepresentations, derogatory and exclusionary language, and other denigrating behaviors that disproportionately affect already-vulnerable and marginalized communities (Bender et al., 2021; Dodge et al., 2021; Sheng et al., 2021b). These harms are forms of "social bias," a subjective and normative term we broadly use to refer to disparate treatment or outcomes between social groups that arise from historical and structural power asymmetries, which we define and discuss in Section 2.111Unless otherwise specified, our use of ”bias” refers to social bias, defined in Definition 7. Though LLMs often reflect existing biases, they can amplify these biases too; in either case, the automated reproduction of injustice can reinforce systems of inequity (Benjamin, 2020). From negative sentiment and toxicity directed towards some social groups, to stereotypical linguistic associations, to lack of recognition of certain language dialects, the presence of biases of LLMs have been well-documented (e.g., Blodgett and O’Connor, 2017; Hutchinson et al., 2020; Mei, Fereidooni, and Caliskan, 2023; Měchura, 2022; Mozafari, Farahbakhsh, and Crespi, 2020; Sap et al., 2019; Sheng et al., 2019).
@@
With the growing recognition of the biases embedded in LLMs has emerged an abundance of works proposing techniques to measure or remove social bias, primarily organized by (1) metrics for bias evaluation, (2) datasets for bias evaluation, and (3) techniques for bias mitigation. In this survey, we categorize, summarize, and discuss each of these areas of research. For each area, we propose an intuitive taxonomy structured around the types of interventions to which a researcher or practitioner has access. Metrics for bias evaluation are organized by the underlying data structure assumed by the metric, which may differ depending on access to the LLM (i.e., can the user access model-assigned token probabilities, or only generated text output?). Datasets are similarly categorized by their structure. Techniques for bias mitigation are organized by the stage of intervention: pre-processing, in-training, intra-processing, and post-processing.
@@
This survey complements existing literature by offering a more extensive and comprehensive examination of bias and fairness in NLP. Surveys of bias and fairness in machine learning, such as Mehrabi et al. (2021) and Suresh and Guttag (2021), offer important broad-stroke frameworks, but are not specific to linguistic tasks or contexts. While previous works within NLP such as Czarnowska, Vyas, and Shah (2021), Kumar et al. (2023b), and Meade, Poole-Dayan, and Reddy (2021) have focused on specific axes of bias evaluation and mitigation, such as extrinsic fairness metrics, empirical validation, and language generation interventions, our work provides increased breadth and depth. Specifically, we offer a comprehensive overview of bias evaluation and mitigation techniques across a wide range of NLP tasks and applications, synthesizing diverse bodies of work to surface unifying themes and overarching challenges. Beyond enumerating techniques, we also examine the limitations of each class of approach, providing insights and recommendations for future work.
@@ 2.1 Preliminaries
A large language model (LLM) $\mathcal{M}$ parameterized by $\theta$ is a model with an autoregressive, autoencoding, or encoder-decoder architecture trained on a corpus of hundreds of millions to trillions of tokens. LLMs encompass pre-trained models.
Autoregressive models include GPT (Radford et al., 2018), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), and GPT-4 (OpenAI, 2023); autoencoding models include BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and XLM-R (Conneau et al., 2020); and encoder-decoder models include BART (Lewis et al., 2020) and T5 (Raffel et al., 2020).
@@
To quantify the performance of an LLM — whether for a downstream task, bias mitigation, or otherwise — an evaluation dataset and metric are typically used. Though benchmark datasets and their associated metrics are often conflated, the evaluation dataset and metric are distinct entities in an evaluation framework, and thus we define a general LLM metric here. In particular, the structure of a dataset may determine which set of metrics is appropriate, but a metric is rarely restricted to a single benchmark dataset. We discuss this relationship in more detail in Sections 3 and 4.
@@ 2.2.1 Social Bias and Fairness
Measuring and mitigating social "bias" to ensure "fairness" in NLP systems has featured prominently in recent literature. Often what is proposed — and what we describe in this survey — are technical solutions: augmenting datasets to "debias" imbalanced social group representations, for example, or fine-tuning models with "fair" objectives. Despite the growing emphasis on addressing these issues, bias and fairness research in LLMs often fails to precisely describe the harms of model behaviors: who is harmed, why the behavior is harmful, and how the harm reflects and reinforces social principles or hierarchies (Blodgett et al., 2020). Many approaches, for instance, assume some implicitly desirable criterion (e.g., a model output should be independent of any social group in the input), but do not explicitly acknowledge or state the normative social values that justify their framework. Others lack consistency in their definitions of bias, or do not seriously engage with the relevant power dynamics that perpetuate the underlying harm (Blodgett et al., 2021). Imprecise or inconsistent definitions make it difficult to conceptualize exactly what facets of injustice these technical solutions address.
@@
Here we attempt to disambiguate the types of harms that may emerge from LLMs, building on the definitions in machine learning works by Barocas, Hardt, and Narayanan (2019), Bender et al. (2021), Blodgett et al. (2020), Crawford (2017), Mehrabi et al. (2021), Suresh and Guttag (2021), and Weidinger et al. (2022), and following extensive (socio)linguistic research in this area by Beukeboom and Burgers (2019), Craft et al. (2020), Loudermilk (2015), Maass (1999), and others. Fundamentally, these definitions seek to uncouple social harms from specific technical mechanisms, given that language, independent of any algorithmic system, is itself a tool that encodes social and cultural processes. Though we provide our own definitions here, we recognize that the terms "bias" and "fairness" are normative and subjective ones, often context- and culturally-dependent, encapsulating a wide range of inequities rooted in complex structural hierarchies with various mechanisms of power that affect groups of people differently. Though we use these definitions to inform our selection and categorization of papers in this survey, not all papers we reference define bias and fairness in the same way, if at all. Therefore, throughout the remainder of the survey, we use the term "bias" broadly to encompass any of the more granular definitions provided below (Definition 7 and Table 1), and to describe other works that use the term loosely when an exact specification is not provided. Note that our use of the terms "debiased" or "unbiased" does not mean that bias has been completely removed, but rather refers to the output of a bias mitigation technique, regardless of that technique’s effectiveness, reflecting language commonly used in prior works. Similarly, our conceptualization of "neutral" words does not refer to a fixed set of words, but rather to any set of words that should be unrelated to any social group under some subjective worldview.
@@
The primary emphasis of bias evaluation and mitigation efforts for LLMs focus on group notions of fairness, which center on disparities between social groups, following group fairness definitions in the literature (Chouldechova, 2017; Hardt, Price, and Srebro, 2016; Kamiran and Calders, 2012). We also discuss individual fairness (Dwork et al., 2012). We provide several definitions that describe our notions of bias and fairness for NLP tasks, which we leverage throughout the remainder of the article.
@@
We highlight that social groups are often socially constructed, a form of classification with delineations that are not static and may be contested (Hanna et al., 2020). The labeling of groups may grant legitimacy to these boundaries, define relational differences between groups, and reinforce social hierarchies and power imbalances, often with very real and material consequences that can segregate, marginalize, and oppress (Beukeboom and Burgers, 2019; Hanna et al., 2020). The harms experienced by each social group vary greatly, due to distinct historical, structural, and institutional forces of injustice that may operate vastly differently for, say, race and gender, and also apply differently across intersectional identities. However, we also emphasize that evaluating and bringing awareness to disparities requires access to social groups. Thus, under the lens of disparity assessment, and following the direction of recent literature in bias evaluation and mitigation for LLMs, we proceed with this notion of social groups. We now define our notions of fairness and bias, in the context of LLMs.
@@
The taxonomy of bias issues synthesizes and consolidates those similarly defined by Barocas, Hardt, and Narayanan (2019), Blodgett et al. (2020), Blodgett (2021), and Crawford (2017). Each form of bias described in Table 1 represents a distinct form of mistreatment, but the harms are not necessarily mutually exclusive nor independent; for instance, representational harms can in turn perpetuate allocational harms. Even though the boundaries between each form of bias may be ambiguous, we highlight Blodgett (2021)’s recommendation that naming specific harms, the different social relationships and histories from which they arise, and the various assumptions made in their conceptualization is important for interrogating the role of NLP technologies in reproducing inequity and injustice. These definitions may also fall under the umbrella of more general notions of safety, which often also lack explicit definitions in research but typically encompass toxic, offensive, or vulgar language (e.g., Kim et al., 2022; Khalatbari et al., 2023; Meade et al., 2023; Ung, Xu, and Boureau, 2022; Xu et al., 2020). Because unsafe language is also intertwined with historical and structural power asymmetries, it provides an alternative categorization of the definitions in Table 1, including in particular derogatory language and toxicity.
We hope that researchers and practitioners can leverage these definitions to describe work in bias mitigation and evaluation with precise language, to identify sociolinguistic harms that exist in the world, to name the specific harms that the work seeks to address, and to recognize the underlying social causes of those harms that the work should take into consideration.
@@
Representational Harms Denigrating and subordinating attitudes towards a social group
Derogatory language Pejorative slurs, insults, or other words or phrases that target and denigrate a social group
e.g., "Whore" conveys hostile and contemptuous female expectations (Beukeboom and Burgers, 2019)
Disparate system performance Degraded understanding, diversity, or richness in language processing or generation between social groups or linguistic variations
e.g., AAE* like "he woke af" is misclassified as not English more often than $\textrm{}^{\dagger}$ equivalents (Blodgett and O’Connor, 2017)
Erasure Omission or invisibility of the language and experiences of a social group
e.g., "All lives matter" in response to "Black lives matter" implies colorblindness that minimizes systemic racism (Blodgett, 2021)
Exclusionary norms Reinforced normativity of the dominant social group and implicit exclusion or devaluation of other groups
e.g., "Both genders" excludes non-binary identities (Bender et al., 2021)
Misrepresentation An incomplete or non-representative distribution of the sample population generalized to a social group
@@
Stereotyping Negative, generally immutable abstractions about a labeled social group
e.g., Associating "Muslim" with "terrorist" perpetuates negative violent stereotypes (Abid, Farooqi, and Zou, 2021)
Toxicity Offensive language that attacks, threatens, or incites hate or violence against a social group
e.g., "I hate Latinos" is disrespectful and hateful (Dixon et al., 2018)
Allocational Harms Disparate distribution of resources or opportunities between social groups
Direct discrimination Disparate treatment due explicitly to membership of a social group
e.g., LLM-aided resume screening may preserve hiring inequities (Ferrara, 2023)
Indirect discrimination Disparate treatment despite facially neutral consideration towards social groups, due to proxies or other implicit factors
@@ 2.2.2 Bias in NLP Tasks
Language is closely tied to identity, social relations, and power. Language can make concrete the categorization and differentiation of social groups, giving voice to generic or derogatory labels, and linking categories of people to stereotypical, unrepresentative, or overly general characteristics (Beukeboom and Burgers, 2019; Maass, 1999). Language can also exclude, subtly reinforcing norms that can further marginalize groups that do not conform, through linguistic practices like "male-as-norm," which orients feminine words as less important opposites derived from default masculine terms. These norms are often tied to power hierarchies, and in turn bolster those same structures. Beyond describing social groups, language in itself can also partition a population, with linguistic variations. Linguistic profiling, for instance, can discriminate against speakers of a dialect considered non-standard (Baugh, 2000; Loudermilk, 2015). In fact, the determination of which forms of language are considered standard or correct also reinforces social hierarchies that can justify the inferiority of some groups (Blodgett et al., 2020; Craft et al., 2020). Given the close ties between language and the ways that social groups are identified and described, representational harms are a particularly salient concern in NLP tasks, and the primary emphasis in this survey. Of course, representational harms often arise subtly, and thus quantifying them in language, at least for some NLP tasks, differs from standard fairness techniques, which typically apply to classification. We provide a non-exhaustive list of examples of settings where bias may manifest in unique forms, depending on the task:
@@
  • Machine Translation: Machine translators may default to masculine words in the case of ambiguity, a form of an exclusionary norm, such as translating from English "I am happy" to the masculine French form "je suis heureux(Měchura, 2022). Defaulting instead to the feminine form "je suis heureuse" may be equally non-ideal if it also assumes a gender out of context.
  • Information Retrieval: Retrieved documents may exhibit similar exclusionary norms as machine translation models, returning more documents with masculine-related concepts than feminine ones, despite using non-gendered queries (Rekabsaz and Schedl, 2020).
@@
  • Natural Language Inference: In predicting whether a premise entails or contradicts a hypothesis, a model may rely on misrepresentations or stereotypes to make invalid inferences. For example, a biased model may infer that "the accountant ate a bagel" entails or contradicts "the man ate a bagel" or "the woman ate a bagel," when the relationship should instead be neutral (Dev et al., 2020).
  • Classification: Toxicity detection models misclassify African-American English tweets as negative more often than those written in Standard American English (Mozafari, Farahbakhsh, and Crespi, 2020; Sap et al., 2019).